OpenStudy (anonymous):

WILL MEDAL AND FAN JUST FOR HELP!!!! question:This prism has a volume of 112 cm3. What would the volume of the prism be if each dimension was doubled? (Scale factor is 2.)

OpenStudy (anonymous):

the length is 8 the height is 7 the width is 2 and they are all centimeters

OpenStudy (anonymous):

@RyanL @robtobey @dtan5457

Directrix (directrix):

If two solids are similar, the cube of the scale factor of the two solids is equal to the ratio of the volumes.

Directrix (directrix):

112/x = (1/2)^2 where x is the volume of the enlarged prism. 112/x =1/8 Cross multiply and solve for x. @likebruh